# VerifAI: Verified Generative AI

### v1

@@ Abstract.
Generative AI has made significant strides, yet concerns about the accuracy and reliability of its outputs continue to grow. Such inaccuracies can have serious consequences such as inaccurate decision-making, the spread of false information, privacy violations, legal liabilities, and more. Although efforts to address these risks are underway, including explainable AI and responsible AI practices such as transparency, privacy protection, bias mitigation, and social and environmental responsibility, misinformation caused by generative AI will remain a significant challenge. We propose that verifying the outputs of generative AI from a data management perspective is an emerging issue for generative AI. This involves analyzing the underlying data from multi-modal data lakes, including text files, tables, and knowledge graphs, and assessing its quality and consistency. By doing so, we can establish a stronger foundation for evaluating the outputs of generative AI models. Such an approach can ensure the correctness of generative AI, promote transparency, and enable decision-making with greater confidence. Our vision is to promote the development of verifiable generative AI and contribute to a more trustworthy and responsible use of AI.
@@ 1. Introduction
Regrettably, real-world data is often incomplete, inconsistent, or inaccurate (Fan and Geerts, 2012; Deng et al., 2017). Researchers across different fields have invested significant effort in addressing these issues. For instance, the database community has developed methods for detecting data errors (Abedjan et al., 2016; Tang et al., 2020), while the natural language community has developed techniques for identifying fake news and misinformation (Sharma et al., 2022; Rodríguez and Iglesias, 2019).
The issue of data quality becomes more important in the era of generative AI, where advanced generative models like ChatGPT (Ouyang et al., 2022), Midjourney, and SlidesAI can generate intricate outputs. This technology has immense potential to transform various fields, including relational data synthesis, natural language question answering, image generation for advertising and marketing, and many others.
@@
  • Inaccurate decision-making: Decisions based on bad data can lead to poor outcomes, including financial losses, or harm to individuals.
@@
  • Damage to trust: The use of bad data can erode trust in organizations and their products, leading to long-term reputational damage.
Although the risks discussed above are already prevalent when using low quality data before the era of generative AI, they are significantly amplified with the widespread adoption of generative models. In addition to the many discussions about regulating generative AI, from a technical perspective there have been two main bodies of work on the topic, explainable AI and responsible AI.
@@
Explainable AI (XAI) (Došilović et al., 2018; Linardatos et al., 2021) seeks to provide explanations for the outputs produced by AI systems. While XAI can be a valuable tool for improving the trustworthiness of generative AI, it may not be able to completely eliminate concerns regarding the accuracy and reliability of generated outputs. This is due in part to the fact that generative AI models can be highly complex, which can make it difficult to provide a complete explanation of how a given output was produced.
Responsible AI (Nargesian et al., 2023) involves the ethical and accountable development and use of AI systems, which encompasses a range of practices such as transparency, privacy protection, bias mitigation, and social and environmental responsibility. Although model providers like OpenAI are making efforts to improve the accuracy and responsible development of generative AI models, the issue of flawed generative data is expected to continue to be a significant challenge.
Our Vision. We advocate for the establishment of a more reliable framework to evaluate the accuracy and reliability of generative AI outputs. We propose a data management perspective for verifying generative AI outputs that is based on identifying supporting data from data lakes and reasoning with it in order to assess the quality and consistency of the generated data. By doing so, we can establish a more robust foundation for evaluating the accuracy and reliability of the data generated by generative AI. We believe that a verification approach can complement the aforementioned approaches in order to ensure that systems using generative AI are deployed responsibly, effectively, and in a trustworthy manner. This process can help maximize the potential benefits of generative AI while minimizing the risks of unintended consequences or negative impacts. It is important to clarify that our proposed approach focuses on verifying generative data that has ground truth, as opposed to subjective data (Halevy, 2020).
@@
Despite efforts from model providers (e.g., OpenAI) and retrieval-enhanced methods (Borgeaud et al., 2021; Ahmad et al., 2023) for improving the accuracy of generative AI during the data generation process, the spread of misinformation from generative AI will remain a significant problem. Our proposed post-generation verification approach can also complement these methods by enhancing the overall reliability of generative AI. By combining our verification approach with those existing and rapidly evolving generative methods, we can create a synergistic effect that further improves the accuracy and reliability of generative AI.
Challenges. Generative AI can generate data in various contexts and domains, utilizing the “world knowledge” that the model has learned (Borgeaud et al., 2021). However, when the generated data is dirty or inaccurate, traditional data cleaning (Ilyas and Chu, 2019; Mahdavi et al., 2019; Wang and Tang, 2014) and data integration (Doan et al., 2012) approaches that solely rely on the data at hand may not be sufficient. Furthermore, when generative AI is used in the context of a specific enterprise, we may need to consult data that is specific to the enterprise in order to verify the correctness of the generated data, raising several new challenges.
  • (C1)
    Indexing and searching multi-modal data lakes. Although there have been efforts to manage data lakes with relational data (Deng et al., 2017; Fernandez et al., 2018a, b; Khatiwada et al., 2023) and textual data (Khattab and Zaharia, 2020), indexing multi-modal data lakes at scale and effectively retrieving top-$k$ data instances cross data modalities remains an unsolved problem.
  • (C2)
    Cross-modal data verification. Data matching is a key concept in data integration (Doan et al., 2012) and data cleaning (Fan et al., 2010). However, matching and reasoning across different data modalities are not well-addressed.
  • (C3)
    Trust of heterogeneous datasets in multi-modal data lakes. Evaluating the trustworthiness of web sources for knowledge fusion has been well studied (Dong et al., 2015). However, evaluating the trustworthiness of different datasets in data lakes, particularly when they are not well curated, remains an open problem.
  • (C4)
    Provenance of the verification process. Fully automated verification of generative AI outputs can be challenging. It is important to store the lineage of the end-to-end verification process, in case the retrieved data from data lakes is flawed or incomplete, or the verification process itself makes mistakes. This allows for later human checks or debugging.
Contributions. In this paper, we present VerifAI, a framework for verified generative AI, which offers the following notable contributions:
  1. (1)
    A modularized framework. We propose a modular framework for verifying generative data that is extendable to different types of data sources and generative data. The framework comprises three main components: an Indexer, a Reranker, and a Verifier, as illustrated in Figure 2 (more details can be found in Section 3).
@@
  2. (2)
    Experiments. We show that VerifAI achieves high accuracy for generated tables and text (see Section 4), which demonstrates the feasibility of using multi-modal data lakes to verify generative AI.
  3. (3)
    Open problems. Contribution (1) highlighted the potential of VerifAI in addressing challenges C1 and C2. However, there are still open problems for these challenges, as well as for other challenges such as C3 and C4, which will be discussed in Section 5.
@@ 2. Verified Generative AI: The Problem
Remark. The verification process is specific to the application and may requires additional metadata from the user. For instance, when given a tuple, the verification requirement could be either on the entire tuple or on a specific column e.g., only verifying the attribute incumbent in Figure 1(a). Another example is when we task a generative AI to fill in the blank for the statement ‘The average score of Michael Jordan’s basketball career is blank’ and then only verify the correctness of the generated number.
@@ 3.1. Indexer
  1. (1)
    Task-agnostic: It can support a wide range of tasks and use cases, making it versatile and adaptable to different needs.
  2. (2)
    Support for both content- and semantic-based search: It can handle traditional string similarity search as well as vector-based search, enabling more advanced search and retrieval capabilities.
Based on the aforementioned principles, VerifAI currently consists of two types of indexes.
The above two types of indexes are commonly used for indexing data lakes. For example, Aurum (Fernandez et al., 2018a) utilizes Elasticsearch, while Microsoft Azure Cognitive Search (https://azure.microsoft.com/en-us/products/search) indexes embeddings (or vectors) of textual data. Combining these two approaches can enhance recall and serve as a foundation for indexing data lakes more effectively, as has been studied in RetClean (Ahmad et al., 2023).
Remark. As these indexes are task-agnostic, the retrieved top-$k$ results may have low accuracy. To ensure the retrieval of correct data instances, $k$ is typically set to a large number (e.g., 100 to 1000) (Khattab and Zaharia, 2020; Ahmad et al., 2023). However, in order to achieve good accuracy with a much smaller number of results after reranking, it is desired to perform task-aware reranking over the retrieved results.
Combiner. While different indexes use different techniques (e.g., content- or semantic-based), their retrieved results typically overlap. The Combiner simply combines these retrieved results from multiple indexes and removes duplicates. Due to its simplicity, we did not depict and discuss it in Figure 2.
@@ 3.2. Reranker
  • (text, text) Reranker: Following RetClean (Ahmad et al., 2023), we also adopt ColBERT (Khattab and Zaharia, 2020) in VerifAI, a holistic approach to compare each token of a query and each token of a retrieved text file, resulting in a more precise score.
@@ 3.3. Verifier
In VerifAI, we utilize two types of Verifiers. The first type is a one-size-fits-all model such as ChatGPT. The second type is composed of specific and localized models designed for different tasks, such as the table fact verification model (Gu et al., 2022b, a) for (text, table) verification, and a fine-tuned RoBERTa (Liu et al., 2019) model for (tuple, tuple) verification (Ahmad et al., 2023), both are from our previous work.
While ChatGPT can be used by default for simplicity, there are two main reasons why we support specific and localized models, which were also discussed in RetClean (Ahmad et al., 2023).
  1. (1)
    Data privacy. Many applications (e.g., healthcare, government) contain highly sensitive information. One legitimate concern when using externally hosted models such as ChatGPT is data privacy. The model could potentially learn and retain sensitive information from the data it has seen (sam, [n.d.]).
  2. (2)
    Better accuracy. Specific and local models, when being fine-tuned on specific datasets and tasks, can oftentimes outperform generic models such as ChatGPT, as will be shown later in Section 4.
Remark. When retrieving data for a generated data instance, it’s possible that we may retrieve multiple data instances that either verify or refute the generated data object. This can occur for several reasons, e.g., incorrect data instances being retrieved. Therefore, understanding the trustworthiness of different data sources (Dong et al., 2015) and maintaining the provenance (Zhang et al., 2021) of verification are important.
@@ 4. Preliminary Results
  • Tuples in need of verification. We collected 100 tuples from web tables. For each tuple, we randomly removed a non-key attribute cell value and then asked ChatGPT to infer the missing value by utilizing the given template provided below. If multiple tuples share the same schema, we can handle them together as a batch.
@@
  • Textual claims in need of verification. We perform a controlled study to assess textual claims, employing 1,300 textual claims from the TabFact (Chen et al., 2019) benchmark, which is currently the most advanced benchmark for verifying the credibility of textual hypotheses by utilizing a given table. This benchmark is suitable for our aim of verifying textual claims. Nevertheless, it’s worth mentioning that although TabFact offers a (textual claim, table) pair, our task demands us to search the data lakes for table(s), which is an exceedingly more difficult task.
@@
Due to space limitations, we will not report the effectiveness of reranking in our study. However, it’s worth noting that reranking (text, text) pairs has been proven effective in ColBERT (Khattab and Zaharia, 2020), and reranking (tuple, tuple) pairs has been discussed in RecClean (Ahmad et al., 2023).
@@
  1. (1)
    When the retrieved data instance supports the imputed tuple or claim, the Verifier outputs “true”;
  2. (2)
    When the retrieved table refutes the imputed tuple or claim, the Verifier outputs “false”;
  3. (3)
    When the retrieved data instance can neither support nor refute it, the Verifier outputs “not related”. To compare ChatGPT with PASTA that only offers two different answers: “true” or “false”, in this case, we consider it’s also correct when PASTA outputs “false”.
5. Open Problems and Call to Arms
@@
Open Problems. We have identified important open problems.
  • Cross-Modal Data Discovery. Data discovery is a challenging problem in data preparation (Deng et al., 2017), particularly in data lakes that contain heterogeneous data stored in various formats (Chen et al., 2023), including structured (e.g., tables), semi-structured (e.g., graphs), and unstructured data. Unlike data lakes that contain only relational tables, discovering data from different modalities requires addressing the heterogeneity of the data. To address this issue, a promising direction is to explore cross-modal representation learning, which involves encoding data from different modalities into a homogeneous vector space. This approach can facilitate a unified data discovery process, such as using a semantic-based indexer, as illustrated in Figure 3.
  • Cross-Modal Verification. In addition to textual and relational data, datasets in other modalities, such as knowledge graph entities (or small subgraphs), can contain valuable information for verifying generative AI. As discussed earlier, generic models like ChatGPT may not provide a comprehensive solution for reasoning due to challenges such as privacy and accuracy. Therefore, a promising direction is to develop local models that are specifically trained for certain use cases, such as (text, knowledge graph entity) or (tuple, text). By focusing on specific cases, these local models can provide more accurate and effective solutions for verifying generative AI.
  • Trustworthiness of Data Sources. The accuracy of discovering and verifying data across different modalities in a data lake can be influenced by the quality and reliability of the underlying data sources. Therefore, it is crucial to assess the trustworthiness of different sources accurately to enhance the overall accuracy and reliability of the entire verification process.
  • Provenance Management. It is important to maintain a record of the provenance of data instances or sources used in verification to facilitate further human checking or debugging.
  • Managing Data Generated by Generative AI. Generative AI solution providers, including OpenAI, are continuously collecting prompts and generated data. While this information is valuable for improving generative AI models, it can also be useful for end-users, particularly enterprise users. Therefore, a promising direction is to explore how to manage the (conversational) prompts and data generated by generative AI to enable better prompt engineering and data lineage tracking, similar to ModelDB (Vartak and Madden, 2018) which is used for managing machine learning models. Such a solution would enable end-users to better facilitate effective collaboration between different stakeholders.
Call to Arms. In this paper, we propose a framework called VerifAI that addresses the growing concern about the reliability of generative AI, which is leading to the spread of misinformation at an alarming rate. The modular design of VerifAI enables the verification of generated data using multi-modal data lakes, paving the way for research activities that will produce practical solutions for enhancing the reliability of generative AI. We discuss several promising directions and describe the early successes of using VerifAI to verify common inaccuracies in data generated by generative AI. We urge the database community to take advantage of this opportunity to make significant advances in this field and join us to improve the reliability of generative AI.

### v2

@@ Abstract.
Generative AI has made significant strides, yet concerns about the accuracy and reliability of its outputs continue to grow. Such inaccuracies can have serious consequences such as inaccurate decision-making, the spread of false information, privacy violations, legal liabilities, and more. Although efforts to address these risks are underway, including explainable AI and AI regulation practices such as transparency, privacy protection, bias mitigation, and social and environmental responsibility, misinformation caused by generative AI will remain a significant challenge. We propose that verifying the outputs of generative AI from a data management perspective is an emerging issue for generative AI. This involves analyzing the underlying data from multi-modal data lakes, including text files, tables, and knowledge graphs, and assessing its quality and consistency. By doing so, we can establish a stronger foundation for evaluating the outputs of generative AI models. Such an approach can ensure the correctness of generative AI, promote transparency, and enable decision-making with greater confidence. Our vision is to promote the development of verifiable generative AI and contribute to a more trustworthy and responsible use of AI.
@@ 1. Introduction
Regrettably, real-world data is often incomplete, inconsistent, or inaccurate. Researchers across different fields have invested significant effort in addressing these issues. For instance, the database community has developed methods for detecting data errors (Abedjan et al., 2016), while the natural language community has developed techniques for identifying fake news and misinformation (Sharma et al., 2022).
The issue of data quality becomes more important in the era of generative AI, where advanced generative models like ChatGPT, Midjourney, and SlidesAI can generate intricate outputs. This technology has immense potential to transform various fields, including relational data synthesis, natural language question answering, image generation for advertising and marketing, and many others.
@@
  1. (1)
    Inaccurate decision-making: Decisions based on bad data can lead to poor outcomes, e.g., financial losses.
  2. (2)
@@
  5. (5)
    Damage to trust: The use of bad data can erode trust in organizations, leading to long-term reputational damage.
Although the risks discussed above are already prevalent when using low quality data before the era of generative AI, they are significantly amplified with the widespread adoption of generative models. In addition to the many discussions about regulating generative AI, from a technical perspective there have been two main bodies of work on the topic, explainable AI and AI regulations.
@@
Explainable AI (XAI) (Linardatos et al., 2021) seeks to provide explanations for the outputs produced by AI systems. While XAI can be a valuable tool for improving the trustworthiness of generative AI, it cannot completely eliminate concerns regarding the accuracy and reliability of generated outputs. This is due in part to the fact that generative AI models can be highly complex, which makes it difficult to provide a complete explanation of how a given output was produced.
AI regulations involve the ethical and accountable development and use of AI systems, which encompasses a range of practices such as transparency, privacy protection, bias mitigation, and social and environmental responsibility. Although model providers like OpenAI and Google are making efforts to regulate the development of generative AI models, the issue of flawed generative data is expected to continue to be a significant challenge.
Our Vision. We advocate for the establishment of a more reliable framework to evaluate the accuracy and reliability of generative AI outputs. We propose a data management perspective for verifying generative AI outputs that is based on identifying supporting data from data lakes and reasoning with it in order to assess the quality and consistency of the generated data. We believe that a verification approach can complement the aforementioned approaches in order to ensure that systems using generative AI are deployed responsibly, effectively, and in a trustworthy manner. It is important to clarify that our proposed approach focuses on verifying generative data that has ground truth, as opposed to subjective data (Halevy, 2020).
@@
Despite efforts from model providers (e.g., OpenAI) and retrieval-enhanced methods (et al., 2021) for improving the accuracy of generative AI during the data generation process, the spread of misinformation from generative AI will remain a significant problem. Our proposed post-generation verification approach can also complement these methods by enhancing the overall reliability of generative AI. By combining our verification approach with those existing and rapidly evolving generative methods, we can create a synergistic effect that further improves the accuracy and reliability of generative AI.
Challenges. Generative AI can generate data in various contexts and domains, utilizing the world knowledge that the model has learned (et al., 2021). However, when the generated data is dirty or inaccurate, traditional data cleaning and integration methods that rely on the data at hand may not be sufficient. Moreover, when generative AI is used in the context of a specific enterprise, we may need to consult data that is specific to the enterprise in order to verify the correctness of the generated data, raising several new challenges.
(C1) Indexing and searching multi-modal data lakes. Although there have been efforts to manage data lakes with relational data (Fernandez et al., 2018) and textual data (Khattab and Zaharia, 2020), indexing multi-modal data lakes at scale and effectively retrieving top-$k$ data instances cross data modalities remains an unsolved problem.
(C2) Cross-modal data verification. Data matching is a key concept in data integration and data cleaning (Doan et al., 2012). However, matching and reasoning across different data modalities are not well-addressed.
(C3) Trust of heterogeneous datasets in multi-modal data lakes. Evaluating the trustworthiness of web sources for knowledge fusion has been well studied (Dong et al., 2015). However, evaluating the trustworthiness of different datasets in data lakes, particularly when they are not well curated, remains an open problem.
(C4) Provenance of the verification process. Fully automated verification of generative AI outputs can be challenging. It is important to store the lineage of the end-to-end verification process, in case the retrieved data from data lakes is flawed or incomplete, or the verification process itself makes mistakes. This allows for later human checks or debugging.
Contributions. In this paper, we present VerifAI, a framework for verified generative AI, which offers the following contributions:
(1) A modularized framework. We propose a modular framework for verifying generative data that is extendable to different types of data sources and generative data. The framework comprises three main components: an Indexer, a Reranker, and a Verifier, as illustrated in Figure 2 (more details can be found in Section 3).
@@
(2) Experiments. We show that VerifAI achieves high accuracy for generated tables and text (Section 4), which demonstrates the feasibility of using multi-modal data lakes to verify generative AI.
(3) Open problems. Contribution (1) highlighted the potential of VerifAI in addressing challenges C1 and C2. However, there are still open problems for these challenges, as well as for other challenges such as C3 and C4, which will be discussed in Section 5.
@@ 2. Verified Generative AI: The Problem
Remark. The verification process is specific to the application and requires additional metadata from the user. For instance, when given a tuple, the verification requirement could be either on the entire tuple or on a specific column e.g., only verifying the attribute incumbent in Figure 1(a).
@@ 3.1. Indexer
(1) Task-agnostic: It can support a wide range of tasks and use cases, making it versatile and adaptable to different needs.
(2) Support for both content- and semantic-based search: It can handle traditional string similarity search as well as vector-based search, enabling more advanced search and retrieval capabilities.
Based on the aforementioned principles, VerifAI currently consists of two indexes.
$\bullet$ Elasticsearch (Gormley and Tong, 2015) supports content-based search, where tables or text files are serialized as strings and then indexed by Elasticsearch.
$\bullet$ Meta Faiss (Johnson et al., 2019) is a library for efficient similarity search and clustering of dense vectors. We first use embedding techniques (e.g., tuple-to-vec or text-to-vec using BERT) to convert tuples or chunked text files into vectors, which are then indexed by Faiss.
The above two types of indexes are commonly used for indexing data lakes. For example, Aurum (Fernandez et al., 2018) utilizes Elasticsearch, while Microsoft Azure Cognitive Search (https://azure.microsoft.com/en-us/products/search) indexes embeddings (or vectors) of textual data. Combining these two approaches can enhance recall and serve as a foundation for indexing data lakes more effectively.
Remark. As these indexes are task-agnostic, the retrieved top-$k$ results may have low accuracy. To ensure the retrieval of correct data instances, $k$ is typically set to a large number (e.g., 100 to 1000 (Khattab and Zaharia, 2020)). However, in order to achieve good accuracy with a much smaller number of results after reranking, it is desired to perform task-aware reranking over the retrieved results.
Combiner. While different indexes use different techniques (e.g., content- or semantic-based,), their retrieved results typically overlap. The Combiner simply combines these retrieved results from multiple indexes and removes duplicates. Due to its simplicity, we did not depict and discuss it in Figure 2.
@@ 3.2. Reranker
-- (text, text) Reranker: We adopt ColBERT (Khattab and Zaharia, 2020), a holistic approach to compare each token of a query and each token of a retrieved text file, resulting in a more precise score.
@@ 3.3. Verifier
In VerifAI, we utilize two types of Verifiers. The first type is a one-size-fits-all model such as ChatGPT. The second type is composed of specific and localized models designed for different tasks, such as the table fact verification model (Gu et al., 2022a) for (text, table) verification, and a fine-tuned RoBERTa model for (tuple, tuple) verification.
While ChatGPT can be used by default for simplicity, there are two main reasons why we support specific and localized models.
$(1)$ Data privacy. Many applications (e.g., healthcare, government) contain highly sensitive information. One legitimate concern when using externally hosted models such as ChatGPT is data privacy. The model could potentially learn and retain sensitive information from the data it has seen (sam, [n.d.]).
$(2)$ Better accuracy. Specific and local models, when being fine-tuned on specific datasets and tasks, can oftentimes outperform generic models such as ChatGPT, as will be shown later in Section 4.
Remark. When retrieving data for a generated data instance, it’s possible that we may retrieve multiple data instances that either verify or refute the generated data object. This can occur for several reasons, e.g., incorrect data instances being retrieved. Therefore, understanding the trustworthiness of different data sources and maintaining the provenance of verification are important.
@@ 4. Preliminary Results
$\bullet$ Tuples in need of verification. We collected 100 tuples from web tables. For each tuple, we randomly removed a non-key attribute cell value and then asked ChatGPT to infer the missing value by utilizing the given template provided below. If multiple tuples share the same schema, we can handle them together as a batch.
@@
$\bullet$ Textual claims in need of verification. We perform a controlled study to assess textual claims, employing 1,300 textual claims from the TabFact (Chen et al., 2019) benchmark, which is currently the most advanced benchmark for verifying the credibility of textual hypotheses by utilizing a given table. This benchmark is suitable for our aim of verifying textual claims. Nevertheless, it’s worth mentioning that although TabFact offers a (textual claim, table) pair, our task demands us to search the data lakes for table(s), which is an exceedingly more difficult task.
@@
Due to space limitations, we will not report the effectiveness of reranking in our study. However, it’s worth noting that reranking (text, text) pairs has been proven effective in ColBERT (Khattab and Zaharia, 2020), and reranking (tuple, tuple) pairs has been discussed in RetClean (Ahmad et al., 2023).
@@
(1) When the retrieved data instance supports the imputed tuple or claim, the Verifier outputs “true”;
(2) When the retrieved table refutes the imputed tuple or claim, the Verifier outputs “false”;
(3) When the retrieved data instance can neither support nor refute it, the Verifier outputs “not related”. To compare ChatGPT with PASTA that only offers two different answers: “true” or “false”, in this case, we consider it’s also correct when PASTA outputs “false”.
@@
5. Open Problems
In this paper, we propose a framework called VerifAI that addresses the growing concern about the reliability of generative AI, which is leading to the spread of misinformation at an alarming rate. The modular design of VerifAI enables the verification of generated data using multi-modal data lakes, paving the way for research activities that will produce practical solutions for enhancing the reliability of generative AI. In addition, we have identified important open problems that can make significant advances in this field and improve the reliability of generative AI.
$\bullet$ Cross-Modal Data Discovery. Data discovery is a challenging problem in data preparation (Fernandez et al., 2018), particularly in data lakes that contain heterogeneous data stored in various formats, including structured (e.g., tables), semi-structured (e.g., graphs), and unstructured data. Unlike data lakes that contain only relational tables, discovering data from different modalities requires addressing the heterogeneity of the data. To address this issue, a promising direction is to explore cross-modal representation learning, which involves encoding data from different modalities into a homogeneous vector space. This approach can facilitate a unified data discovery process, such as using a semantic-based indexer, as illustrated in Figure 3.
@@ 5. Open Problems
$\bullet$ Cross-Modal Verification. In addition to textual and relational data, datasets in other modalities, such as knowledge graph entities (or small subgraphs), can contain valuable information for verifying generative AI. As discussed earlier, generic models like ChatGPT may not provide a comprehensive solution for reasoning due to challenges such as privacy and accuracy. Therefore, a promising direction is to develop local models that are specifically trained for certain use cases, such as (text, knowledge graph entity) or (tuple, text). By focusing on specific cases, these local models can provide more accurate and effective solutions for verifying generative AI.
$\bullet$ Trustworthiness of Data Sources. The accuracy of discovering and verifying data across different modalities in a data lake can be influenced by the quality and reliability of the underlying data sources. Therefore, it is crucial to assess the trustworthiness of different sources accurately to enhance the overall accuracy and reliability of the entire verification process.
$\bullet$ Provenance Management. It is important to maintain a record of the provenance of data instances or sources used in verification to facilitate further human checking or debugging.
$\bullet$ Managing Data Generated by Generative AI. Generative AI solution providers, including OpenAI, are continuously collecting prompts and generated data. While this information is valuable for improving generative AI models, it can also be useful for end-users, particularly enterprise users. Therefore, a promising direction is to explore how to manage the (conversational) prompts and data generated by generative AI to enable better prompt engineering and data lineage tracking, similar to ModelDB (Vartak and Madden, 2018) which is used for managing machine learning models.